A manufacturer is interested in the output voltage of a power supply used in a PC. Output voltage is assumed to be normally dPlease show all work and circle final answers for each individual portion. Thank you.
Answer #1

6-0.3 volt ,D=8 Giveni x = 5.1, Hypothesis Hol Vs Hi- els volts el £ 5 volts test statistic cha Z = X tallo sils 0.3/18 o. 15.2 - S 0.3 3118 0.2 O. Lorl 근 = 1.89 P-ralue P-value = 2* P(Z > Z) = 2* P(Z > 1-89 = 2 * 0.029 (- from P-value of o.58 z-tabP-value Paraluc=2* P(ZKZ) =2* P(Z 4-1.89 o.029 c. from Pralue of z-table o. 058 P-value = 58